OpenStudy (anonymous):

If the height and radius of a cone of volume V are doubled, then the volume of the cone is

hartnn (hartnn):

know the formula for volume of cone ?

hartnn (hartnn):

\(V =(1/3) \pi r^2h\) when , r and h are doubled, \(V_{new} =(1/3) \pi (2r)^2(2h) = 4*2 [(1/3)\pi r^2h]\) so, now can you tell how much times is new volume ?